A. Solve the following.

1. If price increases from 25 to 35 (P=30) and quantity demanded falls from 80 to 60 units (Q=70), what is the elasticity of demand?

2. If price decreases from 40 to 30 and quantity demanded rises from 15 to 30 units, what is the elasticity of demand?

3. If income rises 15 percent in a year and the demand for clothing by 20 percent, find the YED.

4. If 5 percent increase in the of iphone, the demand for Huawei phone increases by 20 percent, find the XED.


1
Expert's answer
2021-10-01T12:26:23-0400
"E=\frac{\varDelta Q}{\varDelta p}\times\frac{p}{Q}=-\frac{6}{7}"

"E=\frac{Q_1-Q_2}{p_1-p_2}\times \frac{p_1+p_2}{Q_1+Q_2}=-\frac{7}{3}"

"E=\frac{\varDelta Q}{\varDelta p}=0.75"

"E=\frac{\varDelta Q}{\varDelta p}=0.25"